OpenStudy (anonymous):

cubed root of 24/125

OpenStudy (lgbasallote):

\[\LARGE \frac{\sqrt[3]{24}}{\sqrt[3]{125}}\]

OpenStudy (anonymous):

Cubed root of 24/125 can be rewritten as cubed root of 24/5^3

OpenStudy (anonymous):

Go from there.

OpenStudy (lgbasallote):

hint rewrite 24 as 8 x 3 8 and 125 are perfect cubes

OpenStudy (anonymous):

no 5to 3 power is 125 not cubed root of 125

OpenStudy (anonymous):

24 is not a perfect cubed and neither is 8

OpenStudy (maheshmeghwal9):